OpenStudy (anonymous):

A parabola has an x-intercept (4,0) and vertex (1,-9). What is the equation of this parabola? a y=-x2+8x-16 b y=x2-6x-8 c y=1/2x2+ 1/2x-10 d y=x2-2x-8

OpenStudy (rajee_sam):

easy way to check is, you are given two points on the curve. (4,0) means when x = 4, y=0 and similarly (1,-9) means when x = 1, y=-9. Just plug in the value of x and see if you get the said y value. That will be your function.

OpenStudy (dumbcow):

do you want the long answer for how to do it if its NOT multiple choice

OpenStudy (anonymous):

Yes please, so I know how to figure it out.

OpenStudy (dumbcow):

ok first you need to know vertex form: \[y = a(x-h)^{2} +k\] (h,k) is vertex in this case: h = 1, k=-9 the other point is : x = 4 , y=0 now plug this in and solve for constant "a" \[0 = a(4-1)^{2} -9\] \[9a-9 = 0\] a = 1 \[y = (x-1)^{2} -9\] expand it out to get it in standard form \[y = (x^{2} -2x+1) -9\] \[y = x^{2} -2x-8\]